# Bayesian Multivariable Bidirectional Mendelian Randomization

### v1

@@ Abstract
Mendelian randomization (MR) is a pivotal tool in genetic epidemiology, leveraging genetic variants as instrumental variables to infer causal relationships between modifiable exposures and health outcomes. Traditional MR methods, while powerful, often rest on stringent assumptions such as the absence of feedback loops, which are frequently violated in complex biological systems. In addition, many popular MR approaches focus on only two variables (i.e., one exposure and one outcome) whereas our motivating applications have many variables. In this article, we introduce a novel Bayesian framework for multivariable MR that concurrently addresses unmeasured confounding and feedback loops. Central to our approach is a sparse conditional cyclic graphical model with a sparse error variance-covariance matrix. Two structural priors are employed to enable the modeling and inference of causal relationships as well as latent confounding structures. Our method is designed to operate effectively with summary-level data, facilitating its application in contexts where individual-level data are inaccessible, e.g., due to privacy concerns. It can also account for horizontal pleiotropy. Through extensive simulations and applications to the GTEx and OneK1K data, we demonstrate the superior performance of our approach in recovering biologically plausible causal relationships in the presence of possible feedback loops and unmeasured confounding. The R package that implements the proposed method is available at MR.RGM.
Keywords: Causal Inference, Instrumental Variables, Unmeasured Confounding, Summary-Level Data, Gene Regulatory Networks, Horizontal Pleiotropy.
@@ 1 Introduction
Mendelian randomization (MR) is a causal inference framework using genetic variants as instrumental variables and has revolutionized the fields of genetics and epidemiology. The main principle of MR is rooted in Mendel’s laws of inheritance, which ensure the random allocation of alleles, thereby mitigating confounding and reverse causation that often plague observational studies davey2003 . This natural randomization resembles the design of randomized controlled trials, offering a powerful alternative when such trials are infeasible.
@@
To date, a wide range of MR methods have been developed including classic methods such as inverse-variance weighting (IVW burgess2013mendelian ) for summary-level data and the two-stage least squares basmann1957generalized ; theil1953repeated for individual-level data as well as more modern approaches such as MR-Egger regression bowden2015mendelian , the weighted median estimator bowden2016 , MR-PRESSO verbanck2018 , the weighted mode-based estimator hartwig2017 , and the generalized summary-data-based Mendelian randomization zhu2018 .
@@
There is also a range of software packages for MR. The MendelianRandomization package R-MendelianRandomization provides R implementations of the IVW, MR-Egger, simple and weighted median, and intercept-based estimators. The TwoSampleMR package R-TwoSampleMR ; Hemani2018 enables systematic two-sample MR analyses using summary-level data, offering a suite of MR methods along with data harmonization tools. The OneSampleMR package R-OneSampleMR is designed for analyses using individual-level data, supporting two-stage predictor substitution and two-stage residual inclusion approaches, which are appropriate when genetic instruments and exposures are measured within the same dataset.
@@
Despite the large literature of MR methods and software packages, the focus has been predominantly on the “one exposure and one outcome” setting, under which only the total causal effect can be inferred. Falling short in addressing the complexities of multiple exposures and outcomes, these MR methods cannot differentiate between direct and indirect/mediation effects. Recognizing the complex nature of biological systems, multivariable MR (MVMR) was developed to estimate the direct effects of multiple exposures on an outcome simultaneously sanderson2019 . MVMR accounts for the correlation between exposures, providing a more nuanced understanding of causal pathways. The mrbayes package R-MrBayes provides Bayesian implementations of IVW and MR-Egger for two-sample MR, including multivariable extensions. GRIVET is a recent approach that infers causal relationships among a set of variables in the presence of unmeasured confounding by leveraging directed acyclic graph (DAG) models chen2023discovery . MrDAG zuber2025bayesian is a Bayesian DAG model utilizing genetic variants as instruments.
@@
While these MVMR approaches have enabled richer causal inference by discerning direct and indirect causal effects, their acyclic assumption excludes the possibility of directed cycles or feedback loops, which are prevalent in many biological systems such as gene regulatory networks, metabolic networks, and phenotypic disease networks. To capture potential feedback loops, Spirtes (1995) spirtes1995directed extends DAG models to directed cyclic graph (DCG) models, which have been applied to genomics cai2013inference ; ni2018reciprocal ; ni2018heterogeneous , brain imaging roy2023directed , and electronic health records jin2025directed . However, they all assume there is no unmeasured confounding and hence would draw biased inference when the assumption is violated.
@@
In this paper, we propose a new Bayesian MVMR approach based on non-recursive structural equation models (SEMs) with sparse correlated errors, termed Mendelian randomization with reciprocal graphical model (MR.RGM), which simultaneously addresses unmeasured confounding and feedback loops using genetic variants as instrumental variables. Besides addressing the challenges of unmeasured confounding and feedback loops, our approach has four additional features. First, it does not require individual-level data; only summary-level data would suffice. In many MR studies, the genetic variants are not publicly available, and only summary statistics are published. Our approach is applicable to those common settings. Second, our approach also infers the exact positions of confounding via a graphical spike-and-slab prior, i.e., the identification of which sets of variables are confounded. Third, the Bayesian nature of our approach allows for natural uncertainty quantification and enhanced interpretability by providing credible intervals for causal effects and posterior inclusion probabilities. Finally, we also extend MR.RGM to account for horizontal pleiotropy and weak instruments.
@@
To assess the performance of our approach, we conduct extensive simulation studies with realistic network structures such as scale-free networks and small-world networks, comparing its performance against existing MR methods that are implemented in the popular MR software packages, MendelianRandomization, mrbayes, and OneSampleMR. Furthermore, we apply the poposed method to two real-world datasets, the GTEx V7 skeletal muscle data gtex2017genetic and the OneK1K data onek1k2022 , successfully recovering biologically plausible causal relationships supported by existing literature.
@@ 2.1 MR, Bidirectional MR, and MVMR
Two-stage least squares (2SLS) and inverse-variance weighting (IVW) are the two most commonly used methods to estimate the causal effect $a_{21}$ of exposure $Y_{1}$ on outcome $Y_{2}$. In 2SLS, we first regress $Y_{1}$ on $X_{1}$ and get a fitted $\widehat{Y}_{1}$ and then regress $Y_{2}$ on $\widehat{Y}_{1}$. The coefficient of $\widehat{Y}_{1}$ in the latter regression is the desired causal effect $a_{21}$. Such an approach can be justified by plugging (1) into (2),
$Y_{2}$ $=a_{21}b_{1}X_{1}+a_{21}c_{1}U+a_{21}E_{1}+c_{2}W+E_{2}$ (3)
$=a_{21}\widehat{Y}_{1}+a_{21}c_{1}U+a_{21}E_{1}+c_{2}W+E_{2},$ (4)
@@
When the direction of causality between $Y_{1}$ and $Y_{2}$ is unknown, separately applying MR in both directions is often adopted, provided that an instrument $X_{2}$ is also available for $Y_{2}$, which is known as the bidirectional MR. When the causal effects in both directions are significant, both may be reported simultaneously, indicating bidirectional/reciprocal causal effects. The validity of using 2SLS and IVW in such a scenario can be justified by considering a non-recursive SEM as a generative model,
@@
where $r_{jk}$ is effect of $X_{k}$ on $Y_{j}$ in a simple linear regression for $j,k\in\{1,2\}$. Hence, the causal effect of $Y_{1}$ on $Y_{2}$ is
@@
which coincides with (5) because the extra factor $\frac{1}{1-a_{12}a_{21}}$ due to the coupling in SEM is common to both $r_{21}$ and $r_{11}$ and thus cancels out in the ratio. Similarly, $a_{12}=\frac{r_{12}}{r_{22}}$. Consequently, both 2SLS and IVW are still valid estimation procedures for bidirectional MR.
@@
where $a_{jk}$ is the direct causal effect of $Y_{k}$ on $Y_{j}$, often represented graphically by $Y_{k}\to Y_{j}$ or simply $k\to j$ in a causal graph. Unlike the bivariate case, applying MR or bidirectional MR to $(Y_{1},Y_{2})$ does not estimate the direct causal effect $a_{12}$. Instead, it targets the total causal effect $t_{12}$ of $Y_{2}$ on $Y_{1}$, which consists of both direct and indirect/mediated effects via $Y_{3}$ and can be found by do-calculus (pearl1995causal, ):
@@
Let $\mathbf{Y}=(Y_{1},\cdots,Y_{p})^{T}$ denote $p$ traits, and let $\mathbf{X}=(X_{1},\cdots,X_{k})^{T}$ represent $k$ instrumental variables. In our later applications, the traits are gene expressions, and the instruments are (cis-) single-nucleotide polymorphisms (SNPs) that are significantly correlated with the traits. Let $\mathbf{U}=(U_{1},\cdots,U_{l})^{T}$ denote a set of $l$ covariates. We model the data-generative process by an SEM:
@@
where $\mathbf{A}=(a_{jh})\in\mathbb{R}^{p\times p}$ with $a_{jh}$ being the direct causal effect of trait $h$ on trait $j$, $\mathbf{B}=(b_{jh})\in\mathbb{R}^{p\times k}$ with $b_{jh}$ capturing the effect of instrumental variable $h$ on trait $j$, $\mathbf{C}=(c_{jh})\in\mathbb{R}^{p\times l}$ with $c_{jh}$ representing the effect of covariate $h$ on trait $j$, $\mathbf{W}=(W_{1},\cdots,W_{t})^{T}$ represents $t$ latent confounders, which are assumed to be $\mathbf{W}\sim\text{N}_{t}(0,\mathbf{I}_{t})$, i.e., $t$ independent sources of unmeasured confounding, $\mathbf{D}=(d_{jh})\in\mathbb{R}^{p\times t}$ with $d_{jh}$ being the impact of unobserved confounder $h$ on trait $j$, and $\mathbf{E}\sim\text{N}_{p}(0,\mathbf{\Sigma})$ with diagonal $\mathbf{\Sigma}$ is the independent exogenous errors. We further assume that there are no self-loops, i.e., $\text{diag}(\mathbf{A})=\mathbf{0}$ and that $\mathbf{X}$, $\mathbf{U}$, $\mathbf{W}$, and $\mathbf{E}$ are all mutually independent. Structural zeros are imposed on $\mathbf{B}$ such that $b_{jh}\neq 0$ if and only if $X_{h}$ is the instrument for $Y_{j}$; this will be later relaxed to account for horizontal pleiotropy and weak instruments.
@@
where $\mathbf{W}$ has been integrated out. The covariance matrix $\mathbf{\Sigma^{*}}$ of this new error term is not diagonal and hence correlated. A non-zero entry $\mathbf{\Sigma}^{*}_{jh}\neq 0$ for $j\neq h$ indicates the presence of a latent confounder affecting both $Y_{j}$ and $Y_{h}$ as it must be exist some $s$ such that neither $d_{js}$ nor $d_{hs}$ is zero. This allows us to infer potential latent confounding structures directly from the covariance matrix of the errors, without having to assume a known number of confounders.
@@
This formulation accommodates feedback loops through $\mathbf{A}$ (e.g., if $a_{jh}\neq 0$ and $a_{hj}\neq 0$, then $Y_{j}\rightleftarrows Y_{h}$) and models latent confounding via the error covariance structure. The conditional distribution of $\mathbf{Y}$ given $\mathbf{X}$ and $\mathbf{U}$ can be derived from (2.2):
@@
$\mathbf{S_{xx}}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{x_{i}}\mathbf{x_{i}}^{T},\quad\mathbf{S_{uu}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{u_{i}}\mathbf{u_{i}}^{T},\quad\mathbf{S_{xu}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{x_{i}}\mathbf{u_{i}}^{T}$
These sufficient statistics allow us to make causal inference without requiring access to individual-level data. The distribution (17), based on the sufficient statistics, is given by:
$\begin{split}&p\left(\{\mathbf{y_{i}}\}_{i=1}^{n}|\{\mathbf{x_{i}}\}_{i=1}^{n},\{\mathbf{u_{i}}\}_{i=1}^{n},\mathbf{A,B,C,\Sigma^{*}}\right)\\ =&(2\pi)^{-\frac{np}{2}}.\mathbf{\text{det}(\Sigma^{*})^{-\frac{n}{2}}.|\text{det}(I_{p}-A)|^{n}}.\exp(-\frac{1}{2}n.\mathbf{[\text{tr}(S_{yy}(I_{p}-A)^{T}{\Sigma^{*}}^{-1}(I_{p}-A))}\\ &\quad\quad\quad\quad\quad\quad-2n.\mathbf{\text{tr}(S_{yx}B^{T}{\Sigma^{*}}^{-1}(I_{p}-A))+n.\text{tr}(S_{xx}B^{T}{\Sigma^{*}}^{-1}B)}\\ &\quad\quad\quad\quad\quad\quad-2n.\text{tr}\mathbf{(S_{yu}^{T}C^{T}{\Sigma^{*}}^{-1}(I_{p}-A))-2n.\text{tr}(S_{xu}C^{T}{\Sigma^{*}}^{-1}B)}\\ &\quad\quad\quad\quad\quad\quad+n.\text{tr}\mathbf{(S_{uu}C^{T}{\Sigma^{*}}^{-1}C)]}).\end{split}$
Our goal is to estimate the matrices $\mathbf{A},\mathbf{B},\mathbf{C}$, and $\mathbf{\Sigma^{*}}$, which together capture the causal relationships among traits, instrumental effects, covariate effects, and the confounding effects. To achieve this, we adopt a fully Bayesian framework: we place prior distributions on these parameters and use Markov chain Monte Carlo (MCMC) to sample them from their joint posterior distribution. Using sufficient statistics also improves the scalability with respect to the sample size $n$. The cost of evaluating the likelihood based on raw data is $\mathcal{O}\!\big(n\{p^{2}+p(k+l)\}+p^{3}\big)$ whereas that of sufficient statistics is $\mathcal{O}\!\big(p^{3}+p^{2}(k+r)+p(k+l)^{2}\big)$, which is a big reduction if $n\gg\{p,k,l\}$. In the following sections, we detail the prior specifications and the posterior inference procedure.

### v2

# Multivariable Bidirectional Mendelian Randomization via Bayesian Directed Cyclic Graphical Models with Correlated Errors

@@ Abstract
Mendelian randomization (MR) is a pivotal tool in genetics, genomics, and epidemiology, leveraging genetic variants as instrumental variables to infer causal relationships between exposures and outcomes. Traditional MR methods, while powerful, often rely on stringent assumptions such as the absence of feedback loops, which are frequently violated in complex biological networks. In addition, many popular MR approaches focus on only two variables (i.e., one exposure and one outcome), whereas our motivating applications of gene regulatory networks have many variables. In this article, we introduce a novel Bayesian framework for multivariable MR that concurrently addresses unmeasured confounding and feedback loops. Central to our approach is a sparse conditional cyclic graphical model with a sparse error variance-covariance matrix. Two structural priors are employed to enable the modeling and inference of causal relationships as well as latent confounding structures. Our method is designed to operate effectively with summary-level data, facilitating its application in contexts where individual-level data are inaccessible, e.g., due to privacy concerns. It can also account for horizontal pleiotropy, under which we establish the sufficient identifiability conditions. Through extensive simulations and applications to the GTEx and OneK11K data, we demonstrate the superior performance of our approach in recovering biologically plausible causal relationships in the presence of possible feedback loops and unmeasured confounding. Using posterior samples, we further quantify uncertainty in inferred network motifs by computing their posterior probabilities. The R package MR.RGM that implements the proposed method is available on CRAN (https://cran.r-project.org/package=MR.RGM).
Keywords: Causal Inference, Identifiability, Feedback Loops, Instrumental Variables, Unmeasured Confounding, Summary-Level Data, Gene Regulatory Networks, Horizontal Pleiotropy, Uncertainty Qunatification.
@@ 1 Introduction
Mendelian randomization (MR) is a causal inference framework using genetic variants as instrumental variables and has revolutionized the fields of genetics and epidemiology. The main principle of MR is rooted in Mendel’s laws of inheritance, which ensure the random allocation of alleles, thereby mitigating confounding and reverse causation that often plague observational studies Davey Smith and Ebrahim (2003). This natural randomization resembles the design of randomized controlled trials, offering a powerful alternative when such trials are infeasible.
@@
Classical MR focuses on a single exposure and a single outcome, graphically represented in Figure 1(a). To date, a wide range of MR methods have been developed including classic methods such as inverse-variance weighting (IVW Burgess et al. (2013)) for summary-level data and two-stage least squares Basmann (1957); Theil (1953) for individual-level data, as well as more modern approaches such as MR-Egger regression Bowden et al. (2015), the weighted median estimator Bowden et al. (2016), MR-PRESSO Verbanck et al. (2018), the weighted mode-based estimator Hartwig et al. (2017), and the generalized summary-data-based Mendelian randomization Zhu et al. (2018).
@@
There is also a range of software packages for MR. The MendelianRandomization package Burgess and Yavorska (2024) provides R implementations of the IVW, MR-Egger, simple and weighted median, and intercept-based estimators. The TwoSampleMR package Hemani et al. (2025, 2018) enables systematic two-sample MR analyses using summary-level data, offering a suite of MR methods along with data harmonization tools. The OneSampleMR package Palmer et al. (2023) is designed for analyses using individual-level data, supporting two-stage predictor substitution and two-stage residual inclusion approaches, which are appropriate when genetic instruments and exposures are measured within the same dataset.
@@
Despite the large literature and software packages for MR methods, the focus has been predominantly on the “one exposure and one outcome” setting, where only the total causal effect can be inferred. Falling short in addressing the complexities of multiple exposures and outcomes, these MR methods cannot differentiate between direct and indirect/mediation effects. Bidirectional MR, for example, uses two separate analyses to test each direction (Figure 1(b)). Recognizing the complex nature of biological networks, multivariable MR (MVMR) was developed to estimate the direct effects of multiple exposures on an outcome simultaneously (Figure 1(c)) Sanderson et al. (2019). MVMR accounts for the correlation between exposures, providing a more nuanced understanding of causal pathways. The mrbayes package Uche-Ikonne et al. (2021) provides Bayesian implementations of IVW and MR-Egger for two-sample MR, including multivariable extensions. GRIVET is a recent approach that infers causal relationships among a set of variables in the presence of unmeasured confounding by leveraging directed acyclic graph (DAG) models Chen et al. (2023). MrDAG Zuber et al. (2025) is a Bayesian DAG model utilizing genetic variants as instruments.
@@
While these MVMR approaches have enabled richer causal inference by discerning direct and indirect causal effects, their acyclic assumption excludes the possibility of directed cycles or feedback loops, which are prevalent in many biological systems such as gene regulatory networks, metabolic networks, and phenotypic disease networks. To capture potential feedback loops, Spirtes (1995) Spirtes (1995) extends DAG models to directed cyclic graph (DCG) models, which have been applied to genomics Cai et al. (2013); Ni et al. (2018a, b), brain imaging Roy et al. (2023), and electronic health records Jin et al. (2025). However, they all assume there is no unmeasured confounding and hence would draw biased inference when the assumption is violated.
@@
In this paper, we propose a new Bayesian MVMR approach based on non-recursive structural equation models (SEMs) with sparse correlated errors, termed Mendelian randomization with reciprocal graphical model (MR.RGM), which simultaneously addresses unmeasured confounding and feedback loops using genetic variants as instrumental variables, while modeling all traits jointly without pre-specifying exposure-outcome roles (Figure 1(d)). Besides addressing the challenges of unmeasured confounding and feedback loops, our approach has four additional features. First, like many MR methods, it does not require individual-level data; only summary-level data would suffice. Often, genetic variants are not publicly available, and only summary statistics are published. Our approach is versatile for these common settings. Second, our approach also infers the exact positions of confounding via a graphical spike-and-slab prior, enabling the identification of which sets of variables may be confounded. Third, the Bayesian nature of our approach allows for natural uncertainty quantification and enhanced interpretability by providing credible intervals for causal effects and posterior edge inclusion probabilities; in addition, we provide a NetworkMotif function that computes posterior probabilities for user-specified network motifs (e.g., a feedforward loop) based on posterior samples. Finally, we also extend MR.RGM to account for horizontal pleiotropy and weak instruments; this extension, which we refer to as MR.RGM+, enables more flexible modeling and requires formal identifiability conditions, which we establish as a theoretical contribution. A conceptual comparison of the aforementioned MR frameworks and ours is summarized in Figure 1.
@@
To assess the performance of our approach, we conduct extensive simulation studies with realistic network structures such as scale-free networks and small-world networks, comparing its performance against existing MR methods that are implemented in the popular MR software packages, MendelianRandomization, mrbayes, and OneSampleMR. Furthermore, we apply the poposed method to two real-world datasets, the GTEx V77 skeletal muscle data Consortium (2017) and the OneK11K data Yazar et al. (2022), where we not only recover biologically plausible causal relationships but also report high posterior probabilities for key network motifs supported by existing biological literature.
$X_{1}$$Y_{1}$$Y_{2}$
(a) Classical MR: one exposure ($Y_{1}$) and one outcome ($Y_{2}$) with instrument $X_{1}$. The direction between exposure and outcome is fixed a priori.
$X_{1}$$Y_{1}$$Y_{2}$$X_{2}$$Y_{2}$$Y_{1}$
(b) Bidirectional MR: two separate MR analyses ($Y_{1}\!\to\!Y_{2}$ and $Y_{2}\!\to\!Y_{1}$), each requiring its own instrument.
$Y_{1}$$Y_{2}$$Y_{3}$$Y_{4}$$X_{1}$$X_{2}$
(c) MVMR: exposures ($Y_{1},Y_{2}$) each with its own instrument; both influence $Y_{3}$, and $Y_{2}$ additionally influences $Y_{4}$. Typically, the direction between exposure and outcomes are fixed a priori.
$X_{1}$$X_{2}$$X_{3}$$X_{4}$$Y_{1}$$Y_{2}$$Y_{3}$$Y_{4}$
(d) Propsoed MR.RGM: supports directed cycles, horizontal pleiotropy, and flexible trait–trait interactions with arbitrary directions (no pre-specification of exposure vs outcome is required).
Figure 1: Conceptual comparison of Mendelian randomization frameworks. (a)–(c) require pre-specified exposure and outcome roles. (a) and (c) do not allow feedback loops. (d) the proposed MR.RGM supports feedback loops, while allowing all traits to be modeled simultaneously without pre-specifying causal direction. Horizontal pleiotropy and unmeasured confounding are handled with principled approach as well.
@@ 2.1 Overview of MR, Bidirectional MR, and MVMR
Two-stage least squares (22SLS) and inverse-variance weighting (IVW) are the two most commonly used methods to estimate the causal effect $a_{21}$ of exposure $Y_{1}$ on outcome $Y_{2}$. In 22SLS, we first regress $Y_{1}$ on $X_{1}$ and get a fitted $\widehat{Y}_{1}$ and then regress $Y_{2}$ on $\widehat{Y}_{1}$. The coefficient of $\widehat{Y}_{1}$ in the latter regression is the desired causal effect $a_{21}$. Such an approach can be justified by plugging (1) into (2),
$Y_{2}$ $=a_{21}b_{1}X_{1}+a_{21}c_{1}W+a_{21}E_{1}+c_{2}W+E_{2}$ (3)
$=a_{21}\widehat{Y}_{1}+a_{21}c_{1}W+a_{21}E_{1}+c_{2}W+E_{2},$ (4)
@@
When the direction of causality between $Y_{1}$ and $Y_{2}$ is unknown, separately applying MR in both directions is often adopted, provided that an instrument $X_{2}$ is also available for $Y_{2}$, which is known as the bidirectional MR. When the causal effects in both directions are significant, both may be reported simultaneously, indicating bidirectional/reciprocal causal effects. The validity of using 22SLS and IVW in such a scenario can be justified by considering a non-recursive SEM as a generative model,
@@
where $r_{jk}$ is the effect of $X_{k}$ on $Y_{j}$ in a simple linear regression for $j,k\in\{1,2\}$. Hence, the causal effect of $Y_{1}$ on $Y_{2}$ is
@@
which coincides with (5) because the extra factor $\frac{1}{1-a_{12}a_{21}}$ due to the coupling in SEM is common to both $r_{21}$ and $r_{11}$ and thus cancels out in the ratio. Similarly, $a_{12}=\frac{r_{12}}{r_{22}}$. Consequently, both 22SLS and IVW are still valid estimation procedures for bidirectional MR even though regression and SEM are generally very different statistical models.
@@
where $a_{jk}$ is the direct causal effect of $Y_{k}$ on $Y_{j}$, often represented graphically by $Y_{k}\to Y_{j}$ or simply $k\to j$ in a causal graph. Unlike the bivariate case, applying MR or bidirectional MR to $(Y_{1},Y_{2})$ does not estimate the direct causal effect $a_{12}$. Instead, it targets the total causal effect $t_{12}$ of $Y_{2}$ on $Y_{1}$, which consists of both direct and indirect/mediated effects via $Y_{3}$ and can be found by do-calculus (Pearl, 1995):
@@
Let $\mathbf{Y}=(Y_{1},\cdots,Y_{p})^{T}$ denote $p$ traits, and let $\mathbf{X}=(X_{1},\cdots,X_{k})^{T}$ represent $k$ instrumental variables. In our later applications, the traits are gene expressions, and the instruments are (cis-) single-nucleotide polymorphisms (SNPs) that are significantly correlated with the traits. Let $\mathbf{U}=(U_{1},\cdots,U_{l})^{T}$ denote a set of $l$ covariates, and $\mathbf{W}=(W_{1},\cdots,W_{t})^{T}$ represents $t$ latent confounders, which are assumed to be $\mathbf{W}\sim\text{N}_{t}(0,\mathbf{I}_{t})$, i.e., $t$ independent sources of unmeasured confounding. We model the data-generative process by an SEM:
@@
where $\mathbf{A}=(a_{jh})\in\mathbb{R}^{p\times p}$ with $a_{jh}$ being the direct causal effect of trait $h$ on trait $j$, $\mathbf{B}=(b_{jh})\in\mathbb{R}^{p\times k}$ with $b_{jh}$ capturing the effect of instrumental variable $h$ on trait $j$, $\mathbf{C}=(c_{jh})\in\mathbb{R}^{p\times l}$ with $c_{jh}$ representing the effect of covariate $h$ on trait $j$, $\mathbf{D}=(d_{jh})\in\mathbb{R}^{p\times t}$ with $d_{jh}$ being the impact of unobserved confounder $h$ on trait $j$, and $\mathbf{E}\sim\text{N}_{p}(0,\mathbf{\Sigma})$ with diagonal $\mathbf{\Sigma}$ is the independent exogenous errors. We further assume that there are no self-loops, i.e., $\text{diag}(\mathbf{A})=\mathbf{0}$ and that $\mathbf{X}$, $\mathbf{U}$, $\mathbf{W}$, and $\mathbf{E}$ are all mutually independent. Structural zeros are imposed on $\mathbf{B}$ such that $b_{jh}\neq 0$ if and only if $X_{h}$ is the instrument for $Y_{j}$; this will be later relaxed to account for horizontal pleiotropy and weak instruments.
@@
where the latent confounders $\mathbf{W}$ have been integrated out. The covariance matrix $\mathbf{\Sigma^{*}}$ of this new error term is not diagonal and hence correlated. A non-zero entry $\mathbf{\Sigma}^{*}_{jh}\neq 0$ for $j\neq h$ indicates the presence of a latent confounder affecting both $Y_{j}$ and $Y_{h}$ as it must exist some $s$ such that neither $d_{js}$ nor $d_{hs}$ is zero. This allows us to infer potential latent confounding structures directly from the covariance matrix of the errors, without having to assume a known number of confounders.
@@
This formulation accommodates feedback loops through $\mathbf{A}$ (e.g., if $a_{jh}\neq 0$ and $a_{hj}\neq 0$, then $Y_{j}\rightleftarrows Y_{h}$) and models latent confounding via the error covariance structure $\mathbf{\Sigma^{*}}$. The conditional distribution of $\mathbf{Y}$ given $\mathbf{X}$ and $\mathbf{U}$ can be derived from (2.2):
@@
$\mathbf{S_{xx}}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{x_{i}}\mathbf{x_{i}}^{T},\quad\mathbf{S_{uu}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{u_{i}}\mathbf{u_{i}}^{T},\quad\mathbf{S_{xu}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{x_{i}}\mathbf{u_{i}}^{T}.$
These sufficient statistics allow us to make causal inference without requiring access to individual-level data. The distribution (17), based on the sufficient statistics, is given by (see Supplementary Section-S11 for the derivation):
$\begin{split}&p\left(\{\mathbf{y_{i}}\}_{i=1}^{n}|\{\mathbf{x_{i}}\}_{i=1}^{n},\{\mathbf{u_{i}}\}_{i=1}^{n},\mathbf{A,B,C,\Sigma^{*}}\right)\\ =&(2\pi)^{-\frac{np}{2}}.\mathbf{\text{det}(\Sigma^{*})^{-\frac{n}{2}}.|\text{det}(I_{p}-A)|^{n}}.\exp(-\frac{1}{2}n.\mathbf{[\text{tr}(S_{yy}(I_{p}-A)^{T}{\Sigma^{*}}^{-1}(I_{p}-A))}\\ &\quad\quad\quad\quad\quad\quad-2n.\mathbf{\text{tr}(S_{yx}B^{T}{\Sigma^{*}}^{-1}(I_{p}-A))+n.\text{tr}(S_{xx}B^{T}{\Sigma^{*}}^{-1}B)}\\ &\quad\quad\quad\quad\quad\quad-2n.\text{tr}\mathbf{(S_{yu}C^{T}{\Sigma^{*}}^{-1}(I_{p}-A))+2n.\text{tr}(S_{xu}C^{T}{\Sigma^{*}}^{-1}B)}\\ &\quad\quad\quad\quad\quad\quad+n.\text{tr}\mathbf{(S_{uu}C^{T}{\Sigma^{*}}^{-1}C)]}).\end{split}$
Our goal is to estimate the matrices $\mathbf{A},\mathbf{B},\mathbf{C}$, and $\mathbf{\Sigma^{*}}$, which respectively capture the causal relationships among traits, instrumental effects, covariate effects, and the confounding effects. To achieve this, we adopt a fully Bayesian framework by placing prior distributions on these parameters and using Markov chain Monte Carlo (MCMC) to sample them from their joint posterior distribution. Using sufficient statistics also improves the scalability with respect to the sample size $n$. The cost of evaluating the likelihood based on raw data is $\mathcal{O}\!\big(n\{p^{2}+p(k+l)\}+p^{3}\big)$ whereas that of sufficient statistics is $\mathcal{O}\!\big(p^{3}+p^{2}(k+l)+p(k+l)^{2}\big)$, which is a big reduction if $n\gg\{p,k,l\}$.